OpenStudy (anonymous):

A stack of nickles and dimes is worth $2.55 . Find the number of nickles if the number of dimes is 12 times less than twice the number of nickles.

OpenStudy (precal):

This is a system of equations

OpenStudy (precal):

Let x = nickels and y = dimes

OpenStudy (precal):

nickle is .05 and dime is .10 correct?

OpenStudy (precal):

.05x + .10y=2.55

OpenStudy (precal):

notice all of the money go together on one line

OpenStudy (precal):

Find the number of nickles if the number of dimes is 12 times less than twice the number of nickles dimes is 12 times less than twice the nickles what did we call dimes?

OpenStudy (precal):

dimes is y is is = 12 less than twice nickels sorry about my typo 12 - 2x

OpenStudy (precal):

2x-12 because when you use less than you have to switch the order so your second equation is y=2x-12

OpenStudy (anonymous):

Im still confused, i get messed up when it comes to math

OpenStudy (precal):

so your system of equations is y=2x-12 .05x+.10y=2.55 so solve your system using any method you would like

OpenStudy (precal):

sorry your might want to look up how to set up systems and solve them. gotta go

OpenStudy (anonymous):

See it is simple, Value of Nickel=0.05 Value of Dime=0.10 Let the number of nickel=x Let the number of dime=y Total value of x number or nickel and y number of dime is 2.55 Hence, 0.05x+0.10y=2.55...................(1) It says that number of nickles if the number of dimes is 12 times less than twice the number of nickles. Twice the number of Nickles=2x 12 times less than twice the number of Nickles=2x-12 which is equal to the number of dimes i.e. y Therefore, the equation becomes.. y=2x-12..........(2) Solve both the equations and lemme know if you still need help :)